# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

### v2

@@ 7 Application to the Italian Mortality Data
For each approach introduced in Section 5, Figure 8 shows the pooled estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 levels. The posterior mean from SDGLMC is also shown (in red) for comparison. The Null model yields the highest estimate on average ($3.92\%$), whereas the GLMadj model estimates the lowest percent change on average ($0.21\%$). The models (i)–(iii) assume that the association of interest is constant over time, however we have argued that this assumption is unlikely.
Figure 8: Estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations: comparison of SDGLMC with the approaches described in Section 5.
The only two approaches able to capture time-varying strengths of the association are the Dummy and the Periodic models. They seem to agree with SDGLMC on the temporal evolution of the effect, showing peaks in summer and troughs in winter. The Dummy and the Periodic models, however, have some limitations, since both assume that the effect changes within the year, but is the same across years. This is not consistent with the evidence from the data, as the pandemic outbreak may have altered the strength of the association.
Although the Dummy model allows for abrupt changes in the exposure effect, these are unlikely in practice as pointed out by Bhaskaran et al. (2013) and Peng et al. (2005). Moreover, the results may be too sensitive to the choice of the boundaries of each season, and a definition based on the calendar seasons may result suboptimal.
In terms of spatial variation, we analyze the differences between area-specific and pooled estimates of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. Figure 9 shows the posterior means of those differences under each competing approach. It also shows the posterior mean of the spatial component of the local exposure effect, i.e. $\delta_{1i}$, expressed in terms of percent change of risk. The spatial variability of the exposure effect approaches zero in all cases except for the Null and GLMadj models. However, when considering the associated uncertainty, there is no evidence of spatial variation in the exposure effect for any of the models. This may be related to the relatively small size of the study region and the similarity between Piemonte and Lombardia (e.g., in terms of socio-economic indicators).
Figure 9: Spatial component of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations under each approach.
@@
In Section G of the Supplementary Material we want to understand the consequences of removing the measured confounders (i.e., temperature, relative humidity, and indicators for weekdays, national holidays, and lockdown periods) from the model specification. Therefore, we impose the restriction that the elements of $\bm{\alpha}$ in Equation (6) are zero. The results are insightful, as the estimated exposure effect greatly overlaps the estimate obtained when the measured confounders are included, with differences occurring only in summer. This suggests that SDGLMC can provide accurate estimates even in complex scenarios, and that the confounding effect may be time-varying and may be due to heat waves.
@@ 8 Discussion
Motivated by a real case-study, in this paper we have proposed the spatial dynamic generalized linear model for confounding (SDGLMC) as a solution for two types of bias that may affect the estimation and interpretation of the exposure–outcome relationship: misspecification and confounding bias. Misspecification bias arises from assuming that the exposure effect is constant in time when it is not. When analyzing the health effects of air pollution, Chiogna and Gaetan (2002) and Peng et al. (2005), among others, argue that there is uncertainty about the “true” functional form of the exposure–outcome association. Therefore, assuming a constant effect may be too restrictive and result in biased estimates. We acknowledge this fact and assume that there exists a smooth (but unknown) function of the exposure, $f(\cdot;\cdot)$, describing the association of interest. Confounding bias is due to concurvity. To control for it, we decompose the exposure effect into large- and small-scale components, then we want to recover the local exposure effect. The same objective is pursued by Janes et al. (2007) and Dupont et al. (2022), but, differently from these authors, we assume that the local effect is spacetime-varying. Hence, SDGLMC constitutes an extension of their models. Importantly, another connection exists between SDGLMC and the work of Dupont et al. (2022), as the smooth function of spatial coordinates in their model and the spatio-temporal intercepts in SDGLMC play a similar role.
We have shown that SDGLMC enjoys some advantages. On one hand, it is able to capture any spacetime-varying exposure effects, thus overcoming the misspecification bias. On the other hand, thanks to the first-order Taylor’s expansion, SDGLMC can model the local exposure effect within a flexible framework. To recover the local, short-term association, the outcome is regressed on a detrended exposure. This is useful because detrending improves estimation stability and helps reduce the confounding bias (Janes et al., 2007, Greven et al., 2011).
The effectiveness of the proposed approach has been demonstrated through artificial data analyses. We have shown that it performs better than several competitors. In particular, SDGLMC not only can reduce the bias, but also can capture complex spatio-temporal dynamics of the exposure effect and almost recover its true functional form. The bias is less alleviated when the unmeasured variable is less smooth than the exposure in time, but it is worth noticing that this situation is rather unrealistic in practice. Note that the models assuming a constant association (e.g., GLMadj) could achieve better performances if the true effect is constant, however the shape of the exposure–outcome association is typically unknown. Unless there is strong prior information about this shape, we thus suggest using SDGLMC.
Many previous empirical works assume a constant exposure–outcome association. However, the evidence from the analysis of the Italian mortality data have indicated that this assumption is conceivably incorrect. In fact, the estimated local exposure effect exhibits a seasonal pattern, with a higher percent change of risk observed during the summer months, when the overall PM2.5 level is lower. Similar conclusions were reached in other studies (Peng et al., 2005, 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018). Although this fact is apparently counterintuitive, it has been observed that it might be due to possible interaction effects of PM2.5 concentrations with meteorological variables, such as temperature, and other air pollutants. We have also noted that our result may be related to possible effect modification due to changes in the particulate matter mixture (Franklin et al., 2008, Hao et al., 2023, Masselot et al., 2022). The relative proportions of the PM2.5 components have not been analyzed in this study, so they could be the subject of future research. However, it is important to note that Hao et al. (2023) found a non-linear long-term association between some PM2.5 constituents and mortality.
A reviewer pointed out that if the effect modification is solely due to measured confounders, it can be studied more straightforwardly by including interaction terms between PM2.5 and the confounders in the linear predictor. We explore this possibility in Section G.2 of the Supplementary Material. However, since the estimated effect is significantly different from that in Figure 7 (especially during the summer months and the pandemic), we conclude that the effect modification is due to both measured and unmeasured confounders.
@@
An exception to the seasonal behavior are the years 2020 and 2021, when the exposure effect is not bounded away from zero. This can be justified by the fact that the lockdown measures, such as stay-at-home orders and the use of masks, have likely reduced the exposure to air pollution. A similar result is found in the absence of measured confounders (see Figure G.1 in the Supplementary Material), and is also suggested by recent studies (see, e.g., Hadei et al., 2021, López-Feldman et al., 2021). It is worth recalling that none of the competing approaches in Section 2 were able to capture this behavior. However, separate analyses per year (see Table 3) suggest patterns that are qualitatively similar to our results. We limited this analysis to the GLMadj and JDZ models due to the limited temporal extent. Nevertheless, both models indicate a non-significant exposure effect during the second phase.
@@
Table 3: Posterior means obtained from fitting the GLMadj and JDZ approaches separately for each year. The values are expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The 95% CIs are reported within the square brackets.
@@
To account for the residual spatial autocorrelation in the outcome, possibly due to unmeasured confounding, SDGLMC accounts for the adjacency of areal units. That is, spatial and space-time components of the intercept are included into the linear predictor. Besides, SDGLMC allows to possibly capture the spatial variations of the local exposure effect, which could potentially discover differences due to urbanicity and morphology of the study area, for example. However, in the application to the Italian mortality data, the local exposure effect is homogeneous across space, probably because the study area is relatively small, and Piemonte and Lombardia have similar socio-economic and environmental characteristics.
@@
The decomposition of the exposure effect into two orthogonal components allows Janes et al. (2007) to test for the presence of unmeasured confounders. The authors state that if the overall and local exposure effects are equal, then it is likely that the confounding bias is not present. In our case, this testing procedure is theoretically possible, but it is not straightforward to implement in practice, since specific assumptions on the functional form of $f(\cdot;\bm{\gamma}_{f})$ are required.
@@
It is arguable that, even if SDGLMC could recover the true exposure effect, $f(\cdot;\cdot)$, its functional form still remains unknown. In other words, the complete list of interaction terms cannot be determined. However, we believe that this is not a drawback only of the proposed model. In fact, the list of interaction terms is unknown when fitting any regression model, and it is up to the researcher to choose the most appropriate form for $f(\cdot;\cdot)$. This decision may involve some degree of uncertainty. The proposed approach, instead, allows for more flexibility and should be able to capture the effect of lockdowns as well as interaction effects between the exposure and other variables, such as other pollutants, so that inferential problems due to misspecification are overcome.
@@
Further research could focus on the causal interpretation of the estimated exposure effect, in which case other assumptions are needed (see, e.g., Hernán and Robins, 2020). However, note that the analytical approach pursued here should not be considered as less credible than causal inference methods (Dominici and Zigler, 2017).
@@
In conclusion, SDGLMC is a flexible and effective approach to estimate the exposure–outcome association in the presence of spacetime-varying effects and unmeasured spatio-temporal confounding. The model is able to capture the complex dynamics of the exposure effect, and to control for confounding. The proposed approach is a valuable tool for researchers interested in the health effects of air pollution, and it can be extended to other fields where the exposure–outcome association is of interest.
@@ Acknowledgements
Hersbach et al. (2018) was downloaded from the C3S (2023). INERIS et al. (2022) was downloaded from the CAMS (2023). The results contain modified C3S and CAMS information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The research of C.Z., P.V., and L.I. is funded by the European Union - NextGenerationEU. C.Z.: PRIN PNRR 2022 (Bando 2023) SLIDE - Stochastic Modeling of Compound Events (P2022KZJTZ - CUP D53D23018920001). P.V. and L.I.: research project PRIN2022 CoEnv - Complex Environmental Data and Modeling (2022E3RY23 - CUP D53D23011080006). A.M.S. is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999).
@@ Appendix A Data Collection
  • Mortality data. All-cause mortality data are collected from the Italian National Institute of Statistics (ISTAT, 2023a) for each municipality in the two regions, and are then spatially aggregated to the HD level, which represents the areal unit. The spatial domain comprises ${N}=117$ HDs where $84$ belong to the Lombardia region and $33$ to the Piemonte region. The municipality of Campione d’Italia is excluded from the analysis since it has a small area and the time series contains many zeroes. Since the data refer to five consecutive years, the temporal dimension is ${T}=1826$.
@@ Appendix B Descriptive Statistics
Descriptive statistics on PM2.5 levels in all HDs during the whole study period are reported in Table B.2. The average daily PM2.5 concentration is $18.21\,\mu g/m^{3}$ (standard deviation = $10.36$) over the entire spatial domain, but Lombardia is more polluted than Piemonte. Table B.2 also reports descriptive statistics of temperature and relative humidity. The average daily temperature is $284.91\,K$ ($7.60$), while the average daily relative humidity is $91.76\%$ ($3.59$). These are very similar to the region-specific values (not shown).
@@ Appendix C Posterior Full Conditional Distributions for the Proposed DGLM
We provide the computational details to make inference on the model proposed in Section 3 and following the discussion in Section 4 of the main manuscript. In particular, stacking Equation (7) of the main text over the ${T}$ time instants, we have (see also Chan and Jeliazkov, 2009):

### v3

@@ 7 Application to the Italian Mortality Data
For each approach introduced in Section 5, Figure 6 shows the pooled estimated percent change in risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 levels. For comparison, the posterior mean from SDGLMC is also shown in red. The Null and GLMadj models yield the highest ($3.92\%$) and the lowest ($0.21\%$) estimates on average, respectively. The Null, GLMadj and JDZ models assume that the association of interest is constant over time; however, we argue that this assumption is unlikely to hold (see Section 1).
Figure 6: Estimated percent change in risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 concentrations: comparison of SDGLMC with the approaches described in Section 5. Vertical dashed lines mark calendar years. The horizontal dotted line indicates no effect.
Among the competing approaches, only the Dummy and Periodic models are able to capture time-varying strengths of the association. They seem to agree with SDGLMC on the temporal evolution of the effect, with peaks in summer and troughs in winter. Although the effect estimates change within the year, they are constant across years. This assumption is limiting, as the pandemic outbreak may have altered the strength of the association. Note also that the Dummy model allows for abrupt changes in the exposure effect; however, these are unlikely in practice as pointed out by Bhaskaran et al. (2013) and Peng et al. (2005). Moreover, the resulting estimates may be too sensitive to the choice of the boundaries of each season, and a definition based on the calendar seasons may result suboptimal.
Regarding spatial variation, we compare the differences between area-specific and pooled estimates under each competing approach with the spatial component of the local exposure effect ($\tilde{\bm{\delta}}_{1}$). Figure 7 shows the posterior means of these quantities, expressed as percent change in risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 concentrations. The spatial variability of the exposure effect approaches zero for all models except the Null and GLMadj. However, when considering the associated uncertainty, there is no evidence of spatial variation in the exposure effect for any of the models. This finding may be related to the relatively small size of the study region and the similarity between Piemonte and Lombardia (e.g., in terms of socio-economic indicators).
Figure 7: Posterior means of the spatial component of the percent change in risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 concentrations under each approach.
@@
In Section G of the Supplementary Material, we investigate the consequences of removing the measured confounders (i.e., temperature, relative humidity, and indicator variables for weekdays, national holidays, and lockdown periods) from the model specification. In other words, we impose the restriction that the elements of $\bm{\alpha}$ in Equation (5) are set to zero. The findings are insightful: the estimated exposure effect greatly overlaps the estimate obtained when the measured confounders are included, with differences occurring only during the summer months. Moreover, it suggests that SDGLMC can provide accurate estimates even in complex scenarios, and that the confounding effect may be time-varying and due to heat waves.
@@ 8 Discussion
Motivated by a real case-study, this paper proposes the spatial dynamic generalized linear model for confounding (SDGLMC) to address two key issues in observational studies that may distort inference and interpretation of the exposure–outcome relationship, namely misspecification bias and confounding bias. Thanks to the first-order Taylor’s expansion, the exposure effect is distinguished into large-scale and small-scale components. To recover the small-scale spatio-temporal association, the outcome is regressed on a detrended exposure. This approach is beneficial because detrending improves estimation stability and helps reduce confounding bias (Janes et al., 2007, Greven et al., 2011). Furthermore, the inclusion of spacetime-varying coefficients allows SDGLMC to overcome both types of bias. As shown in Section 6, SDGLMC is able to almost recover the true functional form of the exposure effect. Thus, it should be preferred over models that assume a constant association.
A similar decomposition of the exposure effect is proposed by Janes et al. (2007) and Dupont et al. (2022); however, they assume a constant effect. Another connection exists between SDGLMC and the work by Dupont et al. (2022), as the smooth function of spatial coordinates in their model and the spatio-temporal intercept in SDGLMC serve a similar role.
The estimated local exposure effect exhibits a seasonal pattern, with a higher percent change in risk observed during the summer months, when the PM2.5 concentrations are generally lower. This behavior indicates that a threshold below which the exposure effect is null may not exist (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). Additionally, the behavior suggests that the unmeasured confounders could have an interaction effect with the exposure (affecting $\beta_{1it}$). The seasonal pattern might be due to potential interactions between PM2.5 concentrations and air temperature. As noted in Section 1, increased ventilation and physiological susceptibility to air pollution may explain the higher percent change in risk during summer. Further, interactions with other variables, such as other pollutants, are also possible (e.g., Liu et al., 2023, Ji et al., 2022, Siddika et al., 2019). Similar findings have been reported in other studies (Nawrot et al., 2007, Peng et al., 2005, 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018). Although the higher exposure effect in periods with lower PM2.5 concentrations may be counterintuitive, the interaction mechanisms described above offer a plausible explanation.
Fine particulate matter is a complex mixture of organic (e.g., black carbon and organic carbon) and inorganic compounds (e.g., sulfate, nitrate, and ammonium), and the relative proportion of these components may vary in time (Franklin et al., 2008, Masselot et al., 2022, Vinson et al., 2020). Their precursors are emitted from both natural (e.g., oceans and volcanoes) and anthropogenic sources, such as fossil fuel combustion, industrial processes, and residential heating. Their transportation and transformation into PM2.5 in the atmosphere are mainly influenced by meteorological conditions (Hao et al., 2023, Jing et al., 2020, Kelly and Fussell, 2012). Toxicological studies indicate that the PM constituents have different effects on human health (e.g., Kelly and Fussell, 2012), and some authors suggest that ammonium is probably among the most harmful components (Masselot et al., 2022, Hao et al., 2023). The relative proportions of the PM2.5 components have not been analyzed in this study, so they could be the subject of future research. Notably, Hao et al. (2023) identify a non-linear long-term association between some PM2.5 constituents and mortality.
@@
An exception to the seasonal behavior is observed in 2020 and 2021, when the exposure effect is not bounded away from zero. This is likely attributable to lockdown measures — such as stay-at-home orders and widespread mask usage — which may have reduced exposure to air pollution. A similar result is found in the absence of measured confounders (see Figure G.1 in the Supplementary Material), and is also supported by recent studies (e.g., Hadei et al., 2021, López-Feldman et al., 2021). It is worth recalling that none of the competing approaches in Section 5 are able to capture the deviation from the seasonal behavior.
The literature reveals a lack of consensus regarding the association between air pollution and adverse COVID-19 health outcomes in the general population (e.g., Brunekreef et al., 2021a, Walton et al., 2021). More recently, Bhaskar et al. (2023) review previous studies and report that the majority suggests a positive association. In contrast, our findings indicate that the local exposure effect in 2020 and 2021 is generally not bounded away from zero (except for a few days). Several factors may explain this result. First, our model uses all-cause mortality among the elderly population as the outcome, making it not directly comparable to COVID-19-specific studies. Second, the issuance of the lockdown measures by the Italian government, along with the mandatory use of facial masks, not only effectively curbed the transmission of the SARS-CoV-2 virus (and of other viruses), but also potentially reduced exposure to air pollution. During lockdown periods, people were required to stay at home, resulting in reduced exposure to outdoor air pollutants. Ventilation was also reduced because lockdowns occurred mainly during winter months. Moreover, the use of facial masks may have reduced the inhalation of airborne hazardous particles (Kodros et al., 2021, Zhou et al., 2018).
Although lockdown periods were not issued in the second half of 2021, the local exposure effect remains not bounded away from zero. This could be due to continued preventing behaviors such as mask-wearing and hand sanitization. Another possible explanation is mortality displacement (or harvesting), where the most vulnerable people may have died earlier than expected in the preceding years due to the SARS-CoV-2 virus in 2020-2021, resulting in an apparently null (or even protective) effect of the exposure afterwards (Armstrong et al., 2017, Bhaskaran et al., 2013, Zanobetti and Schwartz, 2008, Zanobetti et al., 2000).
During the third phase, people started to return to their normal activities, resulting in increased exposure to air pollution. Consequently, the local exposure effect has a shape similar to that in the first phase.
Separate analyses conducted for each year suggest patterns that are qualitatively consistent with our findings. Due to the limited temporal scope, this analysis is restricted to the GLMadj and JDZ models. Table 4 shows that both models indicate a non-significant exposure effect during the second phase.
@@
Table 4: Year-specific estimates of the exposure effect from GLMadj and JDZ models, expressed as percent change in risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. Posterior means are shown for each year from 2018 to 2022, with 95% CIs reported in brackets.
@@
Unmeasured confounding may induce spatial and spatio-temporal correlation in the exposure effect and in the residuals. Thus, SDGLMC includes spatial and spatio-temporal components in the intercept that are structured according to the adjacency of the areal units. Besides, SDGLMC accommodates potential spatial variation in the local exposure effect, which may reveal differences related to urbanicity and morphological characteristics of the study area. However, in the application to the Italian mortality data, the local exposure effect appears homogeneous across space. This is likely due to the relatively small size of the study area and the similar socio-economic and environmental profiles of Piemonte and Lombardia.
A reviewer pointed out that if the effect modification is solely due to measured confounders, it can be more straightforwardly examined by including interaction terms between PM2.5 and the confounders in the linear predictor. We explore this possibility in Section G.2 of the Supplementary Material. However, since the estimated effect is substantially different from that shown in Figure 5 — particularly during the summer months and the pandemic — we conclude that the effect modification is likely due to both measured and unmeasured confounders.
@@
Janes et al. (2007) propose a method for testing the presence of unmeasured confounding when the exposure effect is partitioned into two orthogonal components. According to their approach, if the overall and local exposure effects are equal, it is likely that confounding bias is absent. While this testing procedure is theoretically feasible in our case, its practical implementation is challenging as it requires specific assumptions about the functional form of $f(\cdot;\bm{\gamma}_{f})$.
@@
It is arguable that, even if SDGLMC could recover the true exposure effect $f(\cdot;\cdot)$, its exact functional form remains unknown. In other words, the complete set of interaction terms cannot be determined. However, we believe this limitation is not unique to the proposed model. In fact, the list of interaction terms is generally unknown when fitting any regression model, and it is up to the researcher to choose the most appropriate form for $f(\cdot;\cdot)$. This choice inherently involves a degree of subjectivity and uncertainty. By contrast, the proposed approach offers greater flexibility and should be able to capture the effects of lockdowns as well as interaction effects between exposure and other variables, such as other pollutants, thereby mitigating inferential issues due to misspecification.
@@
Further research could explore the causal interpretation of the estimated exposure effect, which would require additional assumptions (e.g., Hernán and Robins, 2020). However, it is important to emphasize that the analytical approach adopted here should not be considered as less credible than causal inference methods (Dominici and Zigler, 2017).
@@
In conclusion, SDGLMC is a flexible and effective approach to estimate the exposure–outcome association in the presence of spacetime-varying effects and unmeasured spatio-temporal confounding. The model is able to capture the complex dynamics of the exposure effect while controlling for confounding. The proposed approach is a valuable tool for researchers interested in the small-scale, spatio-temporal health effects of air pollution and can be extended to other fields where the exposure–outcome association is of interest.
@@ Acknowledgements
The authors are grateful to the Associate Editor and two anonymous reviewers for their constructive comments that helped improve the manuscript. Hersbach et al. (2018) was downloaded from the C3S (2023). INERIS et al. (2022) was downloaded from the CAMS (2023). The results contain modified C3S and CAMS information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The research of C.Z., P.V., and L.I. is funded by the European Union - NextGenerationEU. C.Z.: PRIN PNRR 2022 (Bando 2023) SLIDE - Stochastic Modeling of Compound Events (P2022KZJTZ - CUP D53D23018920001). P.V. and L.I.: research project PRIN2022 CoEnv - Complex Environmental Data and Modeling (2022E3RY23 - CUP D53D23011080006). A.M.S. is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2024-04312).
Code Availability
The code associated to this paper is available at https://github.com/czaccard/SDGLMC.
@@ Appendix A Data Collection
  • Mortality data. All-cause mortality data are collected from the Italian National Institute of Statistics (ISTAT, 2023a) for each municipality in the two regions, and are then spatially aggregated to the HD level, which represents the areal unit. The spatial domain comprises N=117{N}=117 HDs where $84$ belong to the Lombardia region and $33$ to the Piemonte region. The municipality of Campione d’Italia is excluded from the analysis since it has a small area and the time series contains many zeroes. Since the data refer to five consecutive years, the temporal dimension is T=1826{T}=1826.
@@ Appendix B Descriptive Statistics
Descriptive statistics on PM2.5 levels in all HDs during the whole study period are reported in Table B.2. The average daily PM2.5 concentration is $18.21\,\mu g/m^{3}$ (standard deviation = $10.36$) over the entire spatial domain, but Lombardia is more polluted than Piemonte. Table B.2 also reports descriptive statistics of temperature and relative humidity. The average daily temperature is 284.91K284.91\,K ($7.60$), while the average daily relative humidity is $91.76\%$ ($3.59$). These are very similar to the region-specific values (not shown).
@@ Appendix C Posterior Full Conditional Distributions for the Proposed DGLM
We provide the computational details to make inference on the model proposed in Section 3 and following the discussion in Section 4 of the main manuscript. In particular, stacking Equation (6) of the main text over the T{T} time instants, we have (see also Chan and Jeliazkov, 2009):